Calculate the electric field strength from a 1.25x10-8C charge if the test

charge is placed 2.50x10-5m.


1
Expert's answer
2022-05-20T13:19:38-0400
"E=k\frac{q}{r^2}"

"E=9*10^9*\frac{1.25*10^{-8}}{(2.50*10^{-5})^2}=1.8*10^{11}\:\rm N/C"